ThisGuyHandsome:

𝚂𝚘𝚕𝚟𝚒𝚗𝚐 𝚏𝚘𝚛 𝚝𝚑𝚎 𝚟𝚊𝚛𝚒𝚊𝚋𝚕𝚎 "𝚇"; 𝙰𝚕𝚐𝚎𝚋𝚛𝚊 (𝙍𝙚𝙦𝙪𝙚𝙨𝙩𝙞𝙣𝙜 𝙋𝙤𝙨𝙩 𝙇𝙤𝙘𝙠 𝘼𝙛𝙩𝙚𝙧 𝙏𝙝𝙚 𝙏𝙪𝙩𝙤𝙧𝙞𝙖𝙡 𝙞𝙨 𝘾𝙤𝙢𝙥𝙡𝙚𝙩𝙚𝙙, 𝙋𝙡𝙚𝙖𝙨𝙚 𝘿𝙤𝙣❜𝙩 𝙋𝙤𝙨𝙩 𝘼𝙣𝙮𝙩𝙝𝙞𝙣𝙜, 𝘿𝙈 𝙈𝙚 𝙄𝙛 𝙔𝙤𝙪 𝙁𝙚𝙚𝙡 𝙎𝙤𝙢𝙚𝙩𝙝𝙞𝙣𝙜 𝙎𝙝𝙤𝙪𝙡𝙙 𝘽𝙚 𝘼𝙙𝙙𝙚𝙙 𝙏𝙤 𝙏𝙝𝙚 𝙏𝙪𝙩𝙤𝙧𝙞𝙖𝙡.)

ThisGuyHandsome:

𝚆𝚑𝚎𝚗 𝚝𝚛𝚢𝚒𝚗𝚐 𝚝𝚘 𝚜𝚘𝚕𝚟𝚎 𝚏𝚘𝚛 𝚝𝚑𝚎 "𝚇" 𝚟𝚊𝚛𝚒𝚊𝚋𝚕𝚎 𝚒𝚗 𝚊𝚗 𝚎𝚚𝚞𝚊𝚝𝚒𝚘𝚗, 𝚘𝚗𝚕𝚢 𝚊 𝚌𝚘𝚞𝚙𝚕𝚎 𝚘𝚏 𝚜𝚝𝚎𝚙𝚜 𝚊𝚛𝚎 𝚗𝚎𝚎𝚍𝚎𝚍, 𝚜𝚘 𝚒𝚝 𝚜𝚑𝚘𝚞𝚕𝚍𝚗'𝚝 𝚋𝚎 𝚝𝚘𝚘 𝚕𝚘𝚗𝚐 𝚘𝚏 𝚊 𝚙𝚘𝚜𝚝; 𝙀𝙭𝙖𝙢𝙥𝙡𝙚 𝙀𝙦𝙪𝙖𝙩𝙞𝙤𝙣; \[2(3x-7)+4(3x+2)=6(5x+9)+3\] 𝙵𝚒𝚛𝚜𝚝𝚕𝚢, 𝙼𝚞𝚕𝚝𝚒𝚙𝚕𝚢 𝚎𝚟𝚎𝚛𝚢𝚝𝚑𝚒𝚗𝚐 𝚊𝚌𝚌𝚘𝚛𝚍𝚒𝚗𝚐𝚕𝚢 𝚞𝚜𝚒𝚗𝚐 𝚝𝚑𝚎 𝚛𝚞𝚕𝚎 𝚘𝚏 𝚙𝚎𝚖𝚍𝚊𝚜. \[6x-14+12x+8=30x+54+3\] 𝙏𝙝𝙚𝙣 𝙂𝙧𝙤𝙪𝙥 𝙇𝙞𝙠𝙚 𝙏𝙚𝙧𝙢𝙨 𝙏𝙤𝙜𝙚𝙩𝙝𝙚𝙧(𝙑𝙀𝙍𝙔 𝙄𝙈𝙋𝙊𝙍𝙏𝘼𝙉𝙏); \[18x-6=30x+57\] 𝚂𝚞𝚋𝚝𝚛𝚊𝚌𝚝 𝟷𝟾𝚡 𝚏𝚛𝚘𝚖 𝚋𝚘𝚝𝚑 𝚜𝚒𝚍𝚎𝚜 𝚘𝚏 𝚝𝚑𝚎 𝚎𝚚𝚞𝚊𝚝𝚒𝚘𝚗; \[-6=12x+57\] 𝚃𝚑𝚎𝚗 𝚜𝚞𝚋𝚝𝚛𝚊𝚌𝚝 𝟻𝟽 𝚏𝚛𝚘𝚖 𝚋𝚘𝚝𝚑 𝚜𝚒𝚍𝚎𝚜 𝚘𝚏 𝚝𝚑𝚎 𝚎𝚚𝚞𝚊𝚝𝚒𝚘𝚗; \[-63=12x\] 𝙳𝚒𝚟𝚒𝚍𝚎 𝚋𝚘𝚝𝚑 𝚜𝚒𝚍𝚎𝚜 𝚘𝚏 𝚝𝚑𝚎 𝚎𝚚𝚞𝚊𝚝𝚒𝚘𝚗 𝚋𝚢 𝟷𝟸 𝚊𝚗𝚍 𝚜𝚒𝚖𝚙𝚕𝚒𝚏𝚢. \[x=-\frac{ 63 }{ 12 }=-\frac{ 21}{ 4 }\] 𝙁𝙞𝙣𝙖𝙡𝙡𝙮, 𝙮𝙤𝙪 𝙜𝙚𝙩 𝙮𝙤𝙪𝙧 𝙖𝙣𝙨𝙬𝙚𝙧 𝙖𝙨; \[x=-\frac{ 21}{ 4 }\] 𝙄 𝙝𝙤𝙥𝙚 𝙩𝙝𝙞𝙨 𝙜𝙞𝙫𝙚𝙣 𝙞𝙣𝙛𝙤𝙧𝙢𝙖𝙩𝙞𝙤𝙣 𝙝𝙚𝙡𝙥𝙨; 𝘾𝙤𝙣𝙩𝙖𝙘𝙩 𝙢𝙚 𝙞𝙛 𝙣𝙚𝙚𝙙𝙚𝙙. @𝙏𝙝𝙞𝙨𝙂𝙪𝙮𝙃𝙖𝙣𝙙𝙨𝙤𝙢𝙚

Imagine:

@tranquility Could you post lock this as this is my alternate account. Thank you!